OpenStudy (anonymous):

All identities are equations, and all equations are identities. A. True B. False

OpenStudy (ash2326):

Identities are always true whatever be the value of x or any variable. \[x^2=x\times x\] This is an identity Equations are true for only a set of values \[x^2+2x+1=0\] True for only x=-1 Do you understand this?

OpenStudy (anonymous):

kinda

OpenStudy (ash2326):

So what would be the answer?

OpenStudy (anonymous):

true

OpenStudy (ash2326):

Read my post again

OpenStudy (anonymous):

ok only true for x=-1 so false

OpenStudy (ash2326):

Yes, it's false

OpenStudy (anonymous):

thank you!